OpenStudy (anonymous):

Choose the equation of the horizontal line that passes through the point (2, 3)

OpenStudy (anonymous):

y=mx+b y=0*x+3 y=3 lol

OpenStudy (anonymous):

give a list of equations plz?

OpenStudy (anonymous):

dont need it. horiizontal line is just a line of 0 slope aka m=0. From the point given you can tell it passes thru 3 y=mx+b so y=0x+3 y=3

OpenStudy (anonymous):



OpenStudy (anonymous):

oh okay y=mx +b format ok :)

OpenStudy (anonymous):